# A GPU-accelerated Large-scale Simulator for Transportation System Optimization Benchmarking

### v1

@@ Appendix A Datasets for Transportation System Optimization Benchmarking
Map Building. First, based on mosstool, we selected the bounding boxes as shown in Table 7 for each city for to build the map.
Table 7: Geometry bounding boxes of four cities.
@@
Specifically, we first use code like https://github.com/tsinghua-fib-lab/mosstool/blob/main/examples/map_osm2geojson.py to convert OSM data within the bounding boxes to GeoJSON format, and secondly use code like https://github.com/tsinghua-fib-lab/mosstool/blob/main/examples/build_map.py to build maps from GeoJSON format data. The statistics of the maps of the four cities are shown in Table 8
Table 8: Statistics of the four maps.
@@
Beijing 25945 11953
Shanghai 14837 6270
Paris 14411 6588
New York 19046 8339
Realistic OD Matrix Generation. In order to generate realistic travel demands of the four cities, we perform OD matrix generation based on a diffusion model that has been pre-trained in several regions around the world. The diffusion model is also provided in mosstool.
@@
After completing the above steps, we assign the route based on the shortest time to each vehicles and remove those who are unable to reach their destinations. For each city, we scale the number of vehicles and observe the arrival rate of all vehicles, which refers to the rate of vehicles that successfully reached their destinations, to construct datasets with different congestion levels.
The arrival rate of the dataset is determined as the minimum rate between the morning and evening peak periods. Specifically, an arrival rate of $80\%$ is considered congested, $90\%$ is considered normal, and $95\%$ is considered smooth. Based on the above rates, we construct the travel demand datasets under different congestion levels in the four cities, and the relevant statistics are shown in Table 9.
Table 9: # of trips of datasets.
@@
Beijing 350838 439280 571412
Shanghai 348880 436888 612160
Paris 154276 202664 251236
New York 218712 262706 306078
Appendix B The Settings of Transportation System Optimization Scenarios
@@
B.1 Traffic Signal Control.
@@
B.2 Dynamic Lane Assignment within Junctions.
@@
B.3 Tidal Lane Control.
@@
Candidate Roads Identification. For each city, we extract driving roads from OSM of 2019. We match every road in our map to the road network of 2019, there are three aspects to evaluate the matching, the distance between two roads, the distance between the middle point of road in our map and difference highway level between two roads. Any road that cannot be matched with any roads in 2019 is identified as a newly constructed road, and regarded as a candidate road. The spatial and length statistics of candidate roads are shown in Figure 4. We select the 50 roads with the highest number of vehicles from each candidate set as the optimization set for the algorithm.
Figure 4: The spatial and length distribution of all candidate roads.
Table 10: # of candidate roads.
@@
Appendix C Complete Benchmark Results
C.1 Traffic Signal Control.
Table 11: The benchmark results for the traffic signal control scenario with smooth traffic conditions.
@@ Appendix C Complete Benchmark Results
FixedTime 4426.51 109588 3878.57 120449 3749.65 52459 4369.44 67046
MaxPressure 4099.14 122737 3523.45 132028 3477.37 55123 3951.76 77018
FRAP 4741.04 102556 4261.76 113613 3920.69 51301 4725.86 63572
MPLight 4086.96 121620 3474.95 133044 3486.59 55946 3840.82 78125
CoLight 4757.27 102433 4221.91 115244 3922.72 51351 4714.33 63553
Efficient-MPLight 4529.54 110199 4013.16 121301 3953.45 51236 4390.86 69279
Advanced-MPLight 4750.34 102649 3997.82 120529 3786.51 53982 4378.73 70352
Advanced-CoLight 4740.79 103172 4245.06 114927 3945.78 51164 4748.60 63463
PPO 4005.74 124001 3636.27 130485 3485.31 55915 3914.77 77007
Table 12: The benchmark results for the traffic signal control scenario with normal traffic conditions.
@@
FixedTime 4843.05 120049 4324.07 131020 4245.86 60020 4682.07 72725
MaxPressure 4580.41 132055 4045.75 144640 3984.20 64405 4309.91 83927
FRAP 5105.22 112321 4671.52 121896 4404.23 58481 5002.48 68196
MPLight 4790.12 124991 4674.93 122198 3980.07 64921 4196.37 85331
CoLight 5108.88 112672 4640.50 124565 4413.91 58513 4989.17 68184
Efficient-MPLight 5101.14 113272 4480.28 132995 4364.89 60428 5019.92 67303
Advanced-MPLight 5049.47 117568 4603.57 127911 4281.72 61470 5031.15 67322
Advanced-CoLight 5107.23 112778 4661.07 123336 4408.14 58929 5014.66 68292
PPO 4452.05 136630 4143.19 141768 4017.61 64277 4254.92 84411
Table 13: The benchmark results for the traffic signal control scenario with congested traffic conditions.
@@
FixedTime 5274.90 130419 4928.58 145308 4587.84 66608 4923.74 76165
MaxPressure 5106.47 142654 4790.71 158804 4371.45 71874 4596.29 87897
FRAP 5490.86 122061 5205.55 133039 4743.12 64711 5219.06 71448
MPLight 5496.34 121452 4740.95 160551 4378.05 72643 4468.59 90632
CoLight 5484.53 122508 5185.19 135131 4750.70 64463 5214.91 72091
Efficient-MPLight 5375.55 130151 5104.42 142562 4718.07 65795 5225.18 71556
Advanced-MPLight 5358.05 129512 5129.86 138727 4650.76 67593 5227.87 71713
Advanced-CoLight 5481.23 122564 5207.07 133709 4756.00 64517 5222.79 71697
PPO 4975.47 149589 4814.59 155279 4394.63 71491 4535.97 90111
C.2 Dynamic Lane Assignment within Junctions.
Table 14: The benchmark results for the dynamic lane assignment scenario with smooth traffic conditions.
@@
NoChange 4426.43 109588 3878.56 120449 3749.64 52459 4369.47 67046
Random 4435.44 109450 3875.35 120896 3681.61 53474 4344.24 67271
Rule 4340.58 112462 3807.56 122688 3675.52 53244 4310.52 67737
PPO 4345.00 111391 3804.21 122692 3663.37 53629 4309.67 67626
Table 15: The benchmark results for the dynamic lane assignment scenario with normal traffic conditions.
@@
NoChange 4846.70 119890 4322.21 131145 4245.84 60020 4674.08 72870
Random 4839.71 120338 4324.96 131216 4176.70 61810 4636.11 74055
Rule 4761.33 123673 4258.22 133346 4155.11 62366 4615.01 74254
PPO 4769.98 122929 4256.51 133379 4160.89 61792 4614.52 73907
Table 16: The benchmark results for the dynamic lane assignment scenario with congested traffic conditions.
@@
NoChange 5275.00 130419 4928.51 145308 4587.81 66608 4923.79 76165
Random 5294.29 129976 4932.60 144774 4539.27 68410 4907.50 77460
Rule 5221.05 134661 4875.04 148393 4521.80 68795 4863.84 78469
PPO 5240.12 131854 4874.53 148228 4522.36 68413 4854.74 78857
C.3 Tidal Lane Control.
Table 17: The benchmark results for the tidal lane control scenario with smooth traffic conditions.
@@
NoChange 4422.68 109912 3884.78 120156 3723.22 52992 4388.48 66269
Random 4418.86 110150 3887.78 119940 3711.32 53313 4360.21 67295
Rule 4416.80 109997 3870.30 120366 3687.63 53626 4341.27 68113
PPO 4411.00 110161 3857.94 121344 3674.53 53922 4325.07 68775
Table 18: The benchmark results for the tidal lane control scenario with normal traffic conditions.
@@
NoChange 4844.57 120105 4334.80 130604 4224.94 60794 4675.10 72834
Random 4827.75 120778 4338.40 130283 4216.17 60946 4665.83 73335
Rule 4823.40 120901 4313.48 131315 4192.85 61636 4638.03 74284
PPO 4820.48 120936 4304.29 132167 4187.37 61738 4628.47 74756
Table 19: The benchmark results for the tidal lane control scenario with congested traffic conditions.
@@
NoChange 5278.35 130133 4937.88 144546 4575.58 67201 4923.74 76253
Random 5274.19 130456 4937.78 143921 4568.53 67661 4903.82 77236
Rule 5259.98 131304 4908.70 146094 4555.47 68025 4877.93 79096
PPO 5258.72 131709 4909.53 146145 4544.70 68274 4860.92 79731
C.4 Congestion Pricing.
Table 20: The benchmark results for the congestion pricing scenario with smooth traffic conditions.
@@
NoChange 4433.37 109604 3875.91 120444 3743.18 52611 4368.11 66984
Random 4865.59 96231 3969.04 121137 3705.23 53904 4571.66 63670
$\Delta$-toll 4267.05 118077 3630.38 133056 3611.00 54762 4246.56 72188
EBGtoll 5348.10 75078 4230.05 107683 3759.70 52424 4837.93 55155
Table 21: The benchmark results for the congestion pricing scenario with normal traffic conditions.
@@
NoChange 4840.18 120207 4328.34 130621 4239.23 60100 4681.75 72765
Random 5190.37 105640 4422.00 131346 4144.47 62284 4830.86 68976
$\Delta$-toll 4667.77 131747 4096.62 147533 4040.68 65024 4549.26 78182
EBGtoll 5637.30 80476 4637.64 116624 4240.82 60362 5096.60 59154
Table 22: The benchmark results for the congestion pricing scenario with congested traffic conditions.
@@
NoChange 5281.92 129904 4927.87 145154 4588.83 66603 4928.00 76140
Random 5558.59 115184 5027.27 143400 4488.19 69075 5066.64 72916
$\Delta$-toll 5141.39 143518 4767.74 162829 4399.95 72309 4801.01 82602
EBGtoll 5912.51 86945 5182.92 128301 4571.64 66971 5275.42 63052
C.5 Road Planning.
Table 23: The benchmark results for the road planning scenario with smooth traffic conditions.
@@
No-Change 7411.53 137793 5432.11 151689 5997.46 62016 6954.14 92756
Random 7216.43 139070 5325.55 151828 5899.89 62070 7176.53 88947
Rule 7240.73 138795 5352.50 152124 5907.31 62155 6942.72 92463
SA 7272.01 139326 5358.86 151983 5772.82 62990 6921.12 92754
GeneralBO 7102.82 139983 5297.74 152046 5755.44 62478 6808.24 93858
Table 24: The benchmark results for the road planning scenario with normal traffic conditions.
@@
No-Change 8439.44 161722 6699.66 161722 7193.90 76327 7892.50 105533
Random 8304.71 163148 6567.12 181063 7106.19 75839 7967.24 102996
Rule 8235.49 164247 6570.72 181504 7177.88 76176 7956.08 102951
SA 8332.44 163660 6590.66 180954 7154.16 76164 7871.23 105507
GeneralBO 8242.19 164182 6721.78 178790 7161.74 75979 7759.88 106883
Table 25: The benchmark results for the road planning scenario with congested traffic conditions.
@@
No-Change 9684.37 191786 8793.09 216625 8310.09 87786 8624.63 117606
Random 9496.80 195463 8651.96 219199 8230.79 87704 8663.12 115278
Rule 9489.22 195427 8656.58 218214 8250.89 87872 8682.73 114650
SA 9487.05 195918 8600.00 219711 8188.43 88650 8589.73 118036
GeneralBO 9369.14 198446 8508.84 221049 8161.15 88803 8542.19 118579

### v2

@@ Appendix E Datasets for Transportation System Optimization Benchmarking
Map Building. First, based on the toolchain, we selected the bounding boxes as shown in Table A3 for each city for to build the map.
Table A3: Geometry bounding boxes of four cities.
@@
Specifically, we first convert OSM data within the bounding boxes to GeoJSON format, and secondly build maps from GeoJSON format data. The statistics of the maps of the four cities are shown in Table A4.
Table A4: Statistics of the four maps.
@@
Beijing 25,945 11,953
Shanghai 14,837 6,270
Paris 14,411 6,588
New York 19,046 8,339
Realistic OD Matrix Generation. In order to generate realistic travel demands of the four cities, we perform OD matrix generation based on a diffusion model provided in the toolchain that has been pre-trained in several regions around the world.
@@
After completing the above steps, we use A* algorithm to calculate the route based on the shortest time and assign it to each vehicles and remove those who are unable to reach their destinations. For each city, we scale the number of vehicles and observe the arrival rate of all vehicles, which refers to the rate of vehicles that successfully reached their destinations, to construct datasets with different congestion levels.
The arrival rate of the dataset is determined as the minimum rate between the morning and evening peak periods. Specifically, an arrival rate of $80\%$ is considered congested, $90\%$ is considered normal, and $95\%$ is considered smooth. Based on the above rates, we construct the travel demand datasets under different congestion levels in the four cities, and the relevant statistics are shown in Table A5.
Table A5: #trips of the datasets with different congestion level.
@@
Beijing 350,838 439,280 571,412
Shanghai 348,880 436,888 612,160
Paris 154,276 202,664 251,236
New York 218,712 262,706 306,078
Appendix F The Settings of Transportation System Optimization Scenarios
@@
F.1 Traffic Signal Control.
@@
F.2 Dynamic Lane Assignment within Junctions.
@@
F.3 Tidal Lane Control.
@@
Candidate Roads Identification. For each city, we extract driving roads from OSM of 2019. We match every road in our map to the road network of 2019, there are three aspects to evaluate the matching, the distance between two roads, the distance between the middle point of road in our map and difference highway level between two roads. Any road that cannot be matched with any roads in 2019 is identified as a newly constructed road, and regarded as a candidate road. The spatial and length statistics of candidate roads are shown in Figure A4. We select the 50 roads with the highest number of vehicles from each candidate set as the optimization set for the algorithm.
Figure A4: The spatial and length distribution of all candidate roads.
Table A6: # of candidate roads.
@@
Appendix G Complete Benchmark Results
G.1 Traffic Signal Control.
Table A7: The results for the traffic signal control scenario with smooth traffic conditions.
@@ Appendix G Complete Benchmark Results
FixedTime 4,426.5 109,588 3,878.6 120,449 3,749.7 52,459 4,369.4 67,046
MaxPressure 4,099.1 122,737 3,523.4 132,028 3,477.4 55,123 3,951.8 77,018
FRAP 4,741.0 102,556 4,261.8 113,613 3,920.7 51,301 4,725.9 63,572
MPLight 4,087.0 121,620 3,474.9 133,044 3,486.6 55,946 3,840.8 78,125
CoLight 4,757.3 102,433 4,221.9 115,244 3,922.7 51,351 4,714.3 63,553
Efficient-MPLight 4,529.5 110,199 4,013.2 121,301 3,953.4 51,236 4,390.9 69,279
Advanced-MPLight 4,750.3 102,649 3,997.8 120,529 3,786.5 53,982 4,378.7 70,352
Advanced-CoLight 4,740.8 103,172 4,245.1 114,927 3,945.8 51,164 4,748.6 63,463
PPO 4,005.7 124,001 3,636.3 130,485 3,485.3 55,915 3,914.8 77,007
Table A8: The results for the traffic signal control scenario with normal traffic conditions.
@@
FixedTime 4,843.1 120,049 4,324.1 131,020 4,245.9 60,020 4,682.1 72,725
MaxPressure 4,580.4 132,055 4,045.8 144,640 3,984.2 64,405 4,309.9 83,927
FRAP 5,105.2 112,321 4,671.5 121,896 4,404.2 58,481 5,002.5 68,196
MPLight 4,790.1 124,991 4,674.9 122,198 3,980.1 64,921 4,196.4 85,331
CoLight 5,108.9 112,672 4,640.5 124,565 4,413.9 58,513 4,989.2 68,184
Efficient-MPLight 5,101.1 113,272 4,480.3 132,995 4,364.9 60,428 5,019.9 67,303
Advanced-MPLight 5,049.5 117,568 4,603.6 127,911 4,281.7 61,470 5,031.1 67,322
Advanced-CoLight 5,107.2 112,778 4,661.1 123,336 4,408.1 58,929 5,014.7 68,292
PPO 4,452.1 136,630 4,143.2 141,768 4,017.6 64,277 4,254.9 84,411
Table A9: The results for the traffic signal control scenario with congested traffic conditions.
@@
FixedTime 5,274.9 130,419 4,928.6 145,308 4,587.8 66,608 4,923.7 76,165
MaxPressure 5,106.5 142,654 4,790.7 158,804 4,371.4 71,874 4,596.3 87,897
FRAP 5,490.9 122,061 5,205.6 133,039 4,743.1 64,711 5,219.1 71,448
MPLight 5,496.3 121,452 4,740.9 160,551 4,378.1 72,643 4,468.6 90,632
CoLight 5,484.5 122,508 5,185.2 135,131 4,750.7 64,463 5,214.9 72,091
Efficient-MPLight 5,375.6 130,151 5,104.4 142,562 4,718.1 65,795 5,225.2 71,556
Advanced-MPLight 5,358.1 129,512 5,129.9 138,727 4,650.8 67,593 5,227.9 71,713
Advanced-CoLight 5,481.2 122,564 5,207.1 133,709 4,756.0 64,517 5,222.8 71,697
PPO 4,975.5 149,589 4,814.6 155,279 4,394.6 71,491 4,536.0 90,111
G.2 Dynamic Lane Assignment within Junctions.
Table A10: The results for the dynamic lane assignment scenario with smooth traffic conditions.
@@
NoChange 4,426.4 109,588 3,878.6 120,449 3,749.6 52,459 4,369.5 67,046
Random 4,435.4 109,450 3,875.3 120,896 3,681.6 53,474 4,344.2 67,271
Rule 4,340.6 112,462 3,807.6 122,688 3,675.5 53,244 4,310.5 67,737
PPO 4,345.0 111,391 3,804.2 122,692 3,663.4 53,629 4,309.7 67,626
Table A11: The results for the dynamic lane assignment scenario with normal traffic conditions.
@@
NoChange 4,846.7 119,890 4,322.2 131,145 4,245.8 60,020 4,674.1 72,870
Random 4,839.7 120,338 4,325.0 131,216 4,176.7 61,810 4,636.1 74,055
Rule 4,761.3 123,673 4,258.2 133,346 4,155.1 62,366 4,615.0 74,254
PPO 4,770.0 122,929 4,256.5 133,379 4,160.9 61,792 4,614.5 73,907
Table A12: The results for the dynamic lane assignment scenario with congested traffic conditions.
@@
NoChange 5,275.0 130,419 4,928.5 145,308 4,587.8 66,608 4,923.8 76,165
Random 5,294.3 129,976 4,932.6 144,774 4,539.3 68,410 4,907.5 77,460
Rule 5,221.1 134,661 4,875.0 148,393 4,521.8 68,795 4,863.8 78,469
PPO 5,240.1 131,854 4,874.5 148,228 4,522.4 68,413 4,854.7 78,857
G.3 Tidal Lane Control.
Table A13: The results for the tidal lane control scenario with smooth traffic conditions.
@@
NoChange 4,422.7 109,912 3,884.8 120,156 3,723.2 52,992 4,388.5 66,269
Random 4,418.9 110,150 3,887.8 119,940 3,711.3 53,313 4,360.2 67,295
Rule 4,416.8 109,997 3,870.3 120,366 3,687.6 53,626 4,341.3 68,113
PPO 4,411.0 110,161 3,857.9 121,344 3,674.5 53,922 4,325.1 68,775
Table A14: The results for the tidal lane control scenario with normal traffic conditions.
@@
NoChange 4,844.6 120,105 4,334.8 130,604 4,224.9 60,794 4,675.1 72,834
Random 4,827.8 120,778 4,338.4 130,283 4,216.2 60,946 4,665.8 73,335
Rule 4,823.4 120,901 4,313.5 131,315 4,192.9 61,636 4,638.0 74,284
PPO 4,820.5 120,936 4,304.3 132,167 4,187.4 61,738 4,628.5 74,756
Table A15: The results for the tidal lane control scenario with congested traffic conditions.
@@
NoChange 5,278.4 130,133 4,937.9 144,546 4,575.6 67,201 4,923.7 76,253
Random 5,274.2 130,456 4,937.8 143,921 4,568.5 67,661 4,903.8 77,236
Rule 5,260.0 131,304 4,908.7 146,094 4,555.5 68,025 4,877.9 79,096
PPO 5,258.7 131,709 4,909.5 146,145 4,544.7 68,274 4,860.9 79,731
G.4 Congestion Pricing.
Table A16: The results for the congestion pricing scenario with smooth traffic conditions.
@@
NoChange 4,433.4 109,604 3,875.9 120,444 3,743.2 52,611 4,368.1 66,984
Random 4,865.6 96,231 3,969.0 121,137 3,705.2 53,904 4,571.7 63,670
$\Delta$-toll 4,267.1 118,077 3,630.4 133,056 3,611.0 54,762 4,246.6 72,188
EBGtoll 5,348.1 75,078 4,230.1 107,683 3,759.7 52,424 4,837.9 55,155
Table A17: The results for the congestion pricing scenario with normal traffic conditions.
@@
NoChange 4,840.2 120,207 4,328.3 130,621 4,239.2 60,100 4,681.8 72,765
Random 5,190.4 105,640 4,422.0 131,346 4,144.5 62,284 4,830.9 68,976
$\Delta$-toll 4,667.8 131,747 4,096.6 147,533 4,040.7 65,024 4,549.3 78,182
EBGtoll 5,637.3 80,476 4,637.6 116,624 4,240.8 60,362 5,096.6 59,154
Table A18: The results for the congestion pricing scenario with congested traffic conditions.
@@
NoChange 5,281.9 129,904 4,927.9 145,154 4,588.8 66,603 4,928.0 76,140
Random 5,558.6 115,184 5,027.3 143,400 4,488.2 69,075 5,066.6 72,916
$\Delta$-toll 5,141.4 143,518 4,767.7 162,829 4,399.9 72,309 4,801.0 82,602
EBGtoll 5,912.5 86,945 5,182.9 128,301 4,571.6 66,971 5,275.4 63,052
G.5 Road Planning.
Table A19: The results for the road planning scenario with smooth traffic conditions.
@@
No-Change 7,411.5 137,793 5,432.1 151,689 5,997.5 62,016 6,954.1 92,756
Random 7,216.4 139,070 5,325.6 151,828 5,899.9 62,070 7,176.5 88,947
Rule 7,240.7 138,795 5,352.5 152,124 5,907.3 62,155 6,942.7 92,463
SA 7,272.0 139,326 5,358.9 151,983 5,772.8 62,990 6,921.1 92,754
GeneralBO 7,102.8 139,983 5,297.7 152,046 5,755.4 62,478 6,808.2 93,858
Table A20: The results for the road planning scenario with normal traffic conditions.
@@
No-Change 8,439.4 161,722 6,699.7 161,722 7,193.9 7,6327 7,892.5 105,533
Random 8,304.7 163,148 6,567.1 181,063 7,106.2 75,839 7,967.2 102,996
Rule 8,235.5 164,247 6,570.7 181,504 7,177.9 76,176 7,956.1 102,951
SA 8,332.4 163,660 6,590.7 180,954 7,154.2 76,164 7,871.2 105,507
GeneralBO 8,242.2 164,182 6,721.8 178,790 7,161.7 75,979 7,759.9 106,883
Table A21: The results for the road planning scenario with congested traffic conditions.
@@
No-Change 9,684.4 191,786 8,793.1 216,625 8,310.1 87,786 8,624.6 117,606
Random 9,496.8 195,463 8,652.0 219,199 8,230.8 87,704 8,663.1 115,278
Rule 9,489.2 195,427 8,656.6 218,214 8,250.9 87,872 8,682.7 114,650
SA 9,487.0 195,918 8,600.0 219,711 8,188.4 88,650 8,589.7 118,036
GeneralBO 9,369.1 198,446 8,508.8 221,049 8,161.1 88,803 8,542.2 118,579
Appendix H Web Platform Guidance
Since the simulation and optimization of urban transportation systems needs to include map construction, travel demand generation, simulation, and optimization, its requires interested researchers to invest their time in learning and writing the relevant code. In order to help researchers quickly try out the simulator and simulate and optimize urban transportation systems in any region of the world, we build a web platform to support no-code trials. This platform is open for registration and contains mainly the wizard program with documentation for the simulator as shown in Figure A5.
Figure A5: The home page of the web platform. (best view in color)
Upon entering the Wizard page, the platform provides a complete process of building maps, generating trips, simulation, and optimization.
As shown in Figure A6, you can select any rectangular area on the world map or input the bounding box to submit a map building task. The platform is also pre-built with some of the world’s major cities to select from. You can download the map binary file as the simulation input or continue to use the online platform to generate trip by clicking the Select & Continue button.
Figure A6: The map building page of the web platform. (best view in color)
After building or selecting the map, the second step is to generate the travel demands named as trips in the platform as shown in Figure A7. Following the travel demand construction methodology used in the article as mentioned in Section E, we provide a diffusion model based OD matrix generation method named as AIGC. You can customize the area grid density and total number of people used for OD matrix generation, which will affect the time required for execution. You can also customize the departure time distribution through the web GUI. In addition to OD generation, the platform also automatically performs route selection based on minimum elapsed time to generate the correct inputs that can be used directly for simulation. The results can also download for local execution. Once you have finished generating trips, you can click Select & Continue to access the simulation page.
Figure A7: The trip generation page of the web platform. (best view in color)
As shown in Figure A8, you can start the simulation by simply setting the time range to be simulated. Once the simulation is complete, you can click the Show Result button to view the simulation results. The visualization of the simulation results is based on the WebGL online map rendering framework, in which you can view the movement of vehicles, changes in junction traffic signals, etc., as well as observe the automatically constructed road network and its connectivity.
Figure A8: The simulation page of the web platform. (best view in color)
As shown in Figure A9, the online platform also supports two transportation system optimization scenarios, traffic signal control and dynamic lane assignment within junctions, and provides corresponding baseline methods for selection. You can execute it online and get the results of an hour of training based on the corresponding method.
Figure A9: The optimization page of the web platform. (best view in color)
Moreover, we also write the documentation of our whole system including the simulator and the toolchain to guide researchers that are interested at coding to use the simulator. One of the documentation page is shown in Figure A10.
Figure A10: The documentation page of the web platform. (best view in color)